# Physics-informed DeepCT: Sinogram Wavelet Decomposition Meets Masked Diffusion

### v2

@@ I Introduction
Sparse-view X-ray computed tomography (SVCT) is extensively studied for its low-dose and rapid imaging benefits in medical diagnosis and industrial non-destructive testing [1, 2]. However, due to the incomplete of data acquisition of SVCT, serious artifacts are introduced into the reconstructed images, obscuring important internal structures and features[3]. SVCT reconstruction is a challenging inverse problem and improving the reconstruction quality has been a frontier in recent years[4].
@@
Classical iterative reconstruction methods[5] were proposed to improve the image quality though the performance was still poor on highly sparse views. Compressed sensing[6] utilized priors applicable to sparse data such as total variation (TV)[7, 8] and wavelet frame[9], demonstrating a powerful ability to data with sparsity. However, these methods are computationally expensive due to the iterative update steps required and the effect is limited by the parameter sensitivity[10, 11].
@@
Deep learning-based reconstruction models have gained increasing attention for their high computational speed[12] and robustness of parameter[13]. For example, post-processing methods for CT image domain reconstruction include FBPConvNet[14], dense deconvolution networks[11], and residual encoder-decoder convolutional neural networks[15]. Additionally, hybrid domain processing methods have been proposed to reconstruct high-quality images by learning from both the projection domain and the image domain[16, 17], such as hybrid domain neural network[18]. However, the performance of these methods for image reconstruction is still limited.
@@
In recent years, diffusion models have been increasingly employed in SVCT reconstruction owing to their superior capabilities. Xia et al.[19] introduced a patch-based denoising probabilistic diffusion model to improve SVCT reconstruction performance and address large memory requirements. Additionally, Wu et al.[20] proposed an iteratively optimized data scoring model grounded in SGM to achieve high-quality CT reconstruction for ultra-sparse views. Guan et al.[21] introduced a score-based diffusion model that uses a multi-channel strategy in the projection domain to ensuring that the generated information closer to the original data, resulting in a more accurate SVCT reconstruction. Xu et al.[22] used a diffusion model for stage-by-stage optimization in the wavelet domain, which enhanced sparse-view CT image quality. Xia et al.[23] converted the denoising diffusion probability model into a parallel framework to improve the efficiency of the model, and applied it to the reconstruction of breast CT images in dual-domain sparse view. Yang et al.[24] introduced a dual-domain diffusion model for SVCT reconstruction, which includes a sinogram enhancement module and an image refinement module. Although diffusion models have achieved some success in SVCT reconstruction, they still have certain limitations in capturing finer information[25] and rely on large amounts of high-quality data for training[26].
@@
To partially address these challenges, masked diffusion encourages the model to attend to critical regions or structural features, thereby enhancing its ability to reconstruct local details in data-scarce settings. Moreover, introducing perturbations into the input data can enhance the robustness of the model against missing information and improve the generalization performance of the model. Aversa et al.[27] used a layered diffusion model to generate synthetic segmentation masks for high-fidelity diffusion, reducing reliance on labeled data and preserving image detail and structure. Toker et al.[28] used a denoising diffusion probability model to simultaneously generate images and masks, addressing data scarcity in satellite segmentation tasks while ensuring a wide diversity of samples. Konz et al.[29] applied masked diffusion to medical image segmentation, demonstrating its advantages when dealing with complex anatomical structures and reducing the reliance on large amounts of high-quality training data. With the successful application of masked diffusion, it has been gradually introduced into the field of SVCT reconstruction. For example, Tan et al.[30] proposed a score-based multiscale diffusion model. By introducing a regular mask into the SVCT reconstruction framework, this approach effectively improved model performance. Nevertheless, the reliance on fixed masking strategies imposes constraints on the network’s capacity to learn diverse data distributions, thereby limiting the generalization of the model.
@@
II Preliminary
II-A Sparse-view CT Image Reconstruction
@@
where $y$ represents the measured projected data, $A$ is the system matrix determined by the geometry of CT equipment and the scanning protocol used and $\varepsilon$ represents the system error and random noise. SVCT reconstruction is a classical inverse problem [31]. Owing to the incompleteness in the data acquisition process, accurately reconstructing unknown images from limited measurements presents a significant challenge. The mapping from full projection data $y$ to sparse-view projection data $\hat{y}$ is a process of linear transformation. Specifically, for a linear mapping function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, acting on $y$ by the linear operator $P(\wedge)$, is obtained and expressed as:
@@
[…]
@@
II-B Masked Diffusion
Diffusion models have achieved remarkable success in image, audio, and text fields with their high-quality data generation capability. In the forward process, the model gradually transforms the initial data distribution into a gaussian distribution by progressively adding noise[32, 33, 34, 35]. In the reverse process, the model learns a reverse denoising process to gradually remove the noise, ultimately achieving accurate recovery of the original data. The research on diffusion model mainly focuses on the following aspects: denoising diffusion probabilistic models (DDPMs) [26], score-based generative models (SGMs)[36], and stochastic differential equations(SDEs)[37].
@@ II-B Masked Diffusion
While existing works have leveraged masked diffusion in diverse tasks such as image synthesis[38, 39, 27], image editing[40, 41, 42], image restoration[43, 44], image segmentation[45, 46]. Most approaches focus on natural images and overlook the unique challenges of projection data in medical imaging. In particular, conventional masking strategies fail to simulate the complex patterns of missing or corrupted sinogram data encountered in SVCT reconstruction. To address this gap, by introducing random occlusion in the projection domain, our method enhances the model’s ability to generalize from finite and degenerate inputs.
III Method
III-A Motivation
In medical imaging, reconstructing high-quality images from sparse-view projections is vital for accurate diagnosis. Although traditional deep learning models achieve satisfactory performance when trained and tested within closed, homogeneous datasets. Their generalization ability degrades significantly when deployed on out-of-distribution (OOD) or clinically diverse data. Abdar et al. [47] and Huang et al. [48] emphasize the risk of overfitting to training distributions and the consequent drop in diagnostic reliability. However, current literature rarely explores effective strategies to explicitly enhance robustness to distribution shifts in the projection domain, especially in the context of SVCT. Most approaches focus on either improving reconstruction architectures or increasing data diversity through simple augmentation, failing to systematically address the sensitivity of models to the variability in sampling patterns and acquisition conditions.
@@ III-A Motivation
Proof: Let $y_{M}$ represent the sinogram after random masking $m$. The perturbed data $\tilde{y}$ can be expressed as the sum of the sinogram and its corresponding mask, such that $\tilde{y}=y+y_{M}$. The mean of the perturbed data $\tilde{y}$, denoted as $\tilde{\mu}$, can be formulated as follows:
$\tilde{\mu}=\frac{1}{n}\Sigma^{n}_{i=1}(y_{i}+y_{M_{i}}).$ (4)
The variance of the data after the application of masking can be represented as follows:
$\sigma^{2}(\tilde{y})=\frac{1}{n}\Sigma^{n}_{i=1}((y_{i}-\mu)+(y_{M_{i}}-\mu_{M}))^{2},$ (5)
where $\sigma^{2}(\tilde{y})$ is the variance of $\tilde{y}$, $\mu$ and $\mu_{M}$ represent the mean of $y$ and $y_{M}$, respectively. Depending on the properties of expectation and variance, Eq. (5) can be rewritten as:
$\begin{split}\sigma^{2}(\tilde{y})=\sigma^{2}(y)+\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{M_{i}}-\mu_{M})\\ +\frac{1}{n}\Sigma^{n}_{i=1}(y_{M_{i}}-\mu_{M})^{2},\end{split}$ (6)
where $\sigma^{2}(y)$ is the variance of $y$. For an arbitrary random mask $m_{i}$, we have $y_{M_{i}}=y_{i}\cdot m_{i}$ and $\mu_{M}=\frac{1}{n}\Sigma^{n}_{j=1}y_{j}\cdot m_{j}$. Extracting the cross terms of Eq. (6), then
$\begin{split}\mathcal{F}&=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{M_{i}}-\mu_{M})\\ &=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{i}m_{i}-\frac{1}{n}\Sigma^{n}_{j=1}y_{j}m_{j}).\end{split}$ (7)
$\mathcal{F}$ can be decomposed into the sum of two terms, i.e., $\mathcal{F}=\mathcal{F}_{1}-\mathcal{F}_{2}=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(y_{i}m_{i})-\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)(\frac{1}{n}\Sigma^{n}_{j=1}y_{j}m_{j})$. It implies that $\mathbb{E}[m_{i}]=\frac{a+b}{2}$ since $m_{i}\sim\mathnormal{U}(a,b)$. Let $\mathbb{E}[\cdot]$ stands for expectation. Then we have $\mathbb{E}[\mathcal{F}_{1}]=\frac{1}{n}\Sigma^{n}_{i=1}2(y_{i}-\mu)y_{i}\mathbb{E}[m_{i}]=\frac{a+b}{2}\cdot\frac{2}{n}\Sigma^{n}_{i=1}(y_{i}-\mu)y_{i}$ =0 since $\Sigma^{n}_{i=1}(y_{i}-\mu)y_{i}=\sigma^{2}(y)$ is the variance term, the expectation is 00. Moreover, $\mathbb{E}[\mathcal{F}_{2}]=-\frac{2}{n^{2}}\Sigma^{n}_{i=1}\Sigma^{n}_{j=1}(y_{i}-\mu)\cdot y_{j}\cdot\frac{a+b}{2}=0$ since $\Sigma^{n}_{i=1}(y_{i}-\mu)=0$. Hence, $\mathbb{E}[\mathcal{F}]=0$.
Since $\mathbb{E}[\frac{1}{n}\Sigma^{n}_{i=1}(y_{M_{i}}-\mu_{M})^{2}]\geq 0$, it means that $\mathbb{E}[\sigma^{2}(\tilde{x})]\geq\sigma^{2}(x)$. $\square$
The derivation shows that perturbing the data increases data variance, thereby expanding the distribution from which the model learns. This heuristic improvement in generalization arises from exposing the model to a wider range of data. Notably, the benefit of increased variance hinges on the compatibility between the perturbed data distribution and the target test distribution. Thus, our variance-based formulation should be viewed as a heuristic perspective. Our subsequent experiments (Section IV) empirically confirm that this controlled stochasticity enhances generalization across SVCT reconstruction tasks.
To further enhance image reconstruction fidelity, we leverage the high-frequency components extracted by wavelet transforms, which are crucial for preserving fine structures such as edges and textures. During training, we introduce a randomized perturbation strategy applied specifically to high-frequency wavelet subbands. This enhances the model’s sensitivity to microstructural details and improves its capability to represent multi-scale high-frequency information, thereby boosting detail recovery and robustness under noisey and incomplete data conditions. By integrating high-frequency priors with inherent uncertainty, the model achieves improved generalization while maintaining structural fidelity. Additionally, a two-stage iterative optimization framework is employed to jointly ensure global structural coherence and accurate high-frequency detail reconstruction. This dual-phase strategy balances fidelity and texture synthesis, enabling robust and high-quality reconstructions across diverse imaging scenarios.
@@
III-B Training Process in Sinogram Wavelet Domain
1) Prior Learning for Random Mask Strategy in Sinogram Domain: In the prior learning stage based on the projection domain, we propose a virtual random mask training strategy. Combining the physical characteristics of projection data acquisition and the randomness of embedding masks in the projection domain. The Sinogram Random Mask model (SRM) is constructed through the SDEs to enhance the robustness of the model against distribution offsets. As shown in Fig. 2(a), by applying random masks to the sinograms, SRM introduces random perturbations within the limited data sample space. It effectively augments the diversity of the feature space and sharpens the model focus on a broader spectrum of data distributions. This improves the generalization ability of the model as well as the reliability and accuracy of the reconstruction results.
To tackle challenges such as limited scanning angles, we have devised a specialized random masking strategy. By defining a probability parameter $p\in[0,1]$, this strategy is applied to simulate sparse-view sampling scenarios through stochastic masking of projection data. Specifically, for each training sample: a sparse-view mask is applied with probability $p$. In this case, one masking pattern is randomly selected from a predefined set of sparse-view configurations. With probability $1-p$, no masking is applied, and the full projection data is retained. Given a full-view projection data $y$ and the random mask operator $\hat{m}$, the masked full-view projection data can be expressed as:
@@ III-B Training Process in Sinogram Wavelet Domain
The forward stochastic differential equation (SDE) progressively transforms a complex data distribution into a tractable gaussian distribution by gradually adding noise. Given a diffusion process $\{y(t)\}^{T}_{t=0}$, represented by a continuous time variable $t\in[0,T]$, such that $y(t)\sim p_{t}$, where the sample image data is independently and identically distributed. The general forward process of SDE is expressed as follows[31]:
@@
2) Prior Learning for Random High-frequency in Sinogram Wavelet Domain: In order to improve the detail quality of the reconstructed image, a random high-frequency training strategy based on sinogram wavelets is proposed to obtain the Sinogram wavelet random High-frequency Decomposition model (SHD). In this approach, the projection data is decomposed into frequency bands using orthogonal wavelet basis functions. A dynamic random sampling mechanism is then formulated for high-frequency components, where learnable weights are employed to quantify the salience of each sub-band. By leveraging learnable weights, the significance of each sub-band is assessed, enabling the generation of training data with anisotropic perturbations in the frequency domain. Furthermore, this method clearly distinguishes the high-frequency local details in the projection domain. This forces the network to adaptively focus on the information of key frequency bands.
@@
III-C Cascade Reconstruction of Sinogram and Wavelet

### v3

@@ 1 Introduction
Sparse-view X-ray computed tomography (SVCT) is extensively studied for its low-dose and rapid imaging benefits in medical diagnosis and industrial non-destructive testing [cormack1963representation, hounsfield1973computerized]. However, due to the incomplete of data acquisition of SVCT, serious artifacts are introduced into the reconstructed images, obscuring important internal structures and features[Wang2024A]. SVCT reconstruction is a challenging inverse problem and improving the reconstruction quality has been a frontier in recent years[Li2023Sparse-view].
@@
Classical iterative reconstruction methods[dempster1977maximum] were proposed to improve the image quality though the performance was still poor on highly sparse views. Compressed sensing[donoho2006compressed] utilized priors applicable to sparse data such as total variation (TV)[Sidky2008Image, Yu2009Compressed] and wavelet frame[rioul1991wavelets], demonstrating a powerful ability to data with sparsity. However, these methods are computationally expensive due to the iterative update steps required and the effect is limited by the parameter sensitivity[Han2016Deep, Zhang2018A].
@@
Deep learning-based reconstruction models have gained increasing attention for their high computational speed[Koetzier2023Deep] and robustness of parameter[Zhong2024Impacts]. For example, post-processing methods for CT image domain reconstruction include FBPConvNet[McCann2017Deep], dense deconvolution networks[Zhang2018A], and residual encoder-decoder convolutional neural networks[Chen2017Low-Dose]. Additionally, hybrid domain processing methods have been proposed to reconstruct high-quality images by learning from both the projection domain and the image domain[Zhang2020Artifact, Pan2022Multi-domain], such as hybrid domain neural network[Hu2021Hybrid-Domain]. However, the performance of these methods for image reconstruction is still limited.
@@
In recent years, diffusion models have been increasingly employed in SVCT reconstruction owing to their superior capabilities. Xia et al.[Xia2022Patch-Based] introduced a patch-based denoising probabilistic diffusion model to improve SVCT reconstruction performance and address large memory requirements. Additionally, Wu et al.[Wu2023Data-iterative] proposed an iteratively optimized data scoring model grounded in SGM to achieve high-quality CT reconstruction for ultra-sparse views. Guan et al.[Guan2023Generative] introduced a score-based diffusion model that uses a multi-channel strategy in the projection domain to ensuring that the generated information closer to the original data, resulting in a more accurate SVCT reconstruction. Xu et al.[10403850] used a diffusion model for stage-by-stage optimization in the wavelet domain, which enhanced sparse-view CT image quality. Xia et al.[Xia2023Parallel] converted the denoising diffusion probability model into a parallel framework to improve the efficiency of the model, and applied it to the reconstruction of breast CT images in dual-domain sparse view. Yang et al.[Yang2024A] introduced a dual-domain diffusion model for SVCT reconstruction, which includes a sinogram enhancement module and an image refinement module. Although diffusion models have achieved some success in SVCT reconstruction, they still have certain limitations in capturing finer information[kazerouni2023diffusionmodelsmedicalimage] and rely on large amounts of high-quality data for training[ho2020denoisingdiffusionprobabilisticmodels].
@@
To partially address these challenges, masked diffusion encourages the model to attend to critical regions or structural features, thereby enhancing its ability to reconstruct local details in data-scarce settings. Moreover, introducing perturbations into the input data can enhance the robustness of the model against missing information and improve the generalization performance of the model. Aversa et al.[aversa2024diffinfinite] used a layered diffusion model to generate synthetic segmentation masks for high-fidelity diffusion, reducing reliance on labeled data and preserving image detail and structure. Toker et al.[toker2024satsynthaugmentingimagemaskpairs] used a denoising diffusion probability model to simultaneously generate images and masks, addressing data scarcity in satellite segmentation tasks while ensuring a wide diversity of samples. Konz et al.[konz2024anatomicallycontrollablemedicalimagegeneration] applied masked diffusion to medical image segmentation, demonstrating its advantages when dealing with complex anatomical structures and reducing the reliance on large amounts of high-quality training data. With the successful application of masked diffusion, it has been gradually introduced into the field of SVCT reconstruction. For example, Tan et al.[tan2024msdiff] proposed a score-based multiscale diffusion model. By introducing a regular mask into the SVCT reconstruction framework, this approach effectively improved model performance. Nevertheless, the reliance on fixed masking strategies imposes constraints on the network’s capacity to learn diverse data distributions, thereby limiting the generalization of the model.
@@
2 Preliminary
2.1 Sparse-view CT Image Reconstruction
@@
where $y$ represents the measured projected data, $A$ is the system matrix determined by the geometry of CT equipment and the scanning protocol used and $\varepsilon$ represents the system error and random noise. SVCT reconstruction is a classical inverse problem [song2021solving]. Owing to the incompleteness in the data acquisition process, accurately reconstructing unknown images from limited measurements presents a significant challenge. The mapping from full projection data $y$ to sparse-view projection data $\hat{y}$ is a process of linear transformation. Specifically, for a linear mapping function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, acting on $y$ by the linear operator $P(\wedge)$, is obtained and expressed as:
@@
$y^{\ast}=\underset{y}{\arg\min}\ \frac{1}{2}\lVert P(\wedge)y-\hat{y}\rVert^{2}_{2}+\frac{\nu}{2}R(y),$ (3)
@@
2.2 Masked Diffusion
Diffusion models have achieved remarkable success in image, audio, and text fields with their high-quality data generation capability. In the forward process, the model gradually transforms the initial data distribution into a gaussian distribution by progressively adding noise[acar2024advanced, ozturk2024diffusion, ozbey2023unsupervised, gungor2023adaptive]. In the reverse process, the model learns a reverse denoising process to gradually remove the noise, ultimately achieving accurate recovery of the original data. The research on diffusion model mainly focuses on the following aspects: denoising diffusion probabilistic models (DDPMs) [ho2020denoisingdiffusionprobabilisticmodels], score-based generative models (SGMs)[song2019generative], and stochastic differential equations(SDEs)[song2020score].
@@ 2.2 Masked Diffusion
While existing works have leveraged masked diffusion in diverse tasks such as image synthesis[wang2024flame, gao2023masked, aversa2024diffinfinite], image editing[couairon2022diffedit, wang2023instructedit, zou2024towards], image restoration[pang2024improved, zhu2023denoising], image segmentation[le2024maskdiff, toker2024satsynth]. Most approaches focus on natural images and overlook the unique challenges of projection data in medical imaging. In particular, conventional masking strategies fail to simulate the complex patterns of missing or corrupted sinogram data encountered in SVCT reconstruction. To address this gap, by introducing random occlusion in the projection domain, our method enhances the model’s ability to generalize from finite and degenerate inputs.
3 Method
3.1 Motivation
In medical imaging, reconstructing high-quality images from sparse-view projections is vital for accurate diagnosis. Although traditional deep learning models achieve satisfactory performance when trained and tested within closed datasets. Their generalization ability degrades significantly when deployed on out-of-distribution (OOD) or clinically diverse data. Abdar et al. [ABDAR2021243] and Huang et al. [HUANG2025103334] emphasize the risk of overfitting to training distributions and the consequent drop in diagnostic reliability. However, current literature rarely explores effective strategies to explicitly enhance robustness to distribution shifts in the projection domain, especially in the context of SVCT. Most approaches focus on either improving reconstruction architectures or increasing data diversity through simple augmentation, failing to systematically address the sensitivity of models to the variability in sampling patterns and acquisition conditions.
@@ 3.1 Motivation
See Appendix II for proof.
Proposition 3.2: Let the original data sample space be $y=\{y_{1},y_{2},\dots,y_{n}\}$, and $\tilde{y}=y+m\odot y$ denote the masked sample space, where $m_{i}\sim\mathnormal{U}(0,1)$. The covariance of the masked data satisfies $\tilde{\Sigma}\geq\Sigma$. Consequently, for any direction $\nu\in\mathbb{R}^{d}$, it holds that $\nu^{T}\tilde{\Sigma}\nu\geq\nu^{T}\Sigma\nu$, which may lead to the expansion of the data distribution in the masked data space.
See Appendix III for proof.
Our analysis indicates that perturbed data increase the variance of the dataset, thereby expanding the distribution that the model learns from. This form of uncertainty-driven data perturbation helps enhance the model’s generalization ability. It is worth noting that the benefit of increased variance depends on the compatibility between the perturbed data distribution and the target test distribution. Therefore, our variance-based formulation should be regarded as a heuristic perspective. Subsequent experiments demonstrate that this controlled randomness improves the generalization capability in the SVCT reconstruction task.
Additionally, a two-stage iterative optimization framework is employed to jointly ensure global structural coherence and accurate high-frequency detail reconstruction. This dual-phase strategy balances fidelity and texture synthesis, enabling robust and high-quality reconstructions across diverse imaging scenarios.
@@
3.2 Training Process in Sinogram Wavelet Domain
3.2.1 Prior Learning for Random Mask Strategy in Sinogram Domain
In the prior learning stage based on the projection domain, we propose a virtual random mask training strategy. Combining the physical characteristics of projection data acquisition and the randomness of embedding masks in the projection domain. The Sinogram Random Mask model (SRM) is constructed through the SDEs to enhance the robustness of the model against distribution offsets. As shown in Fig. 2(a), by applying random masks to the sinograms, SRM introduces random perturbations within the limited data sample space. It effectively augments the diversity of the feature space and sharpens the model focus on a broader spectrum of data distributions. This improves the generalization ability of the model as well as the reliability and accuracy of the reconstruction results.
By defining a probability parameter $p\in[0,1]$, this strategy is applied to simulate sparse-view sampling scenarios through stochastic masking of projection data. Specifically, for each training sample: a sparse-view mask is applied with probability $p$. In this case, one masking pattern is randomly selected from a predefined set of sparse-view configurations. Given a full-view projection data $y$ and the random mask operator $\hat{m}$, the masked full-view projection data can be expressed as:
@@ 3.2.1 Prior Learning for Random Mask Strategy in Sinogram Domain
The forward stochastic differential equation (SDE) progressively transforms a complex data distribution into a tractable gaussian distribution by gradually adding noise. Given a diffusion process $\{y(t)\}^{T}_{t=0}$, represented by a continuous time variable $t\in[0,T]$, such that $y(t)\sim p_{t}$, where the sample image data is independently and identically distributed. The general forward process of SDE is expressed as follows[song2021solving]:
@@
3.2.2 Prior Learning for Random High-frequency in Sinogram Wavelet Domain
In order to improve the detail quality of the reconstructed image, a random high-frequency training strategy based on sinogram wavelets is proposed to obtain the Sinogram wavelet random High-frequency Decomposition model (SHD). In this approach, the projection data is decomposed into frequency bands using orthogonal wavelet basis functions. A dynamic random sampling mechanism is then formulated for high-frequency components, where learnable weights are employed to quantify the salience of each sub-band. By leveraging learnable weights, the significance of each sub-band is assessed, enabling the generation of training data with anisotropic perturbations in the frequency domain. Furthermore, this method clearly distinguishes the high-frequency local details in the projection domain. This forces the network to adaptively focus on the information of key frequency bands.
@@ 3.2.2 Prior Learning for Random High-frequency in Sinogram Wavelet Domain
3.3 Cascade Reconstruction of Sinogram and Wavelet